OpenStudy (anonymous):

how to find the missing denominator/numerator in multiplying rational expressions

OpenStudy (agent47):

can u post an example?

OpenStudy (anonymous):

cn u give an example of what you are asking?

OpenStudy (anonymous):

_______/x+2 . x+2/x-3 = x+3

OpenStudy (anonymous):

so how can i find the numerator??

OpenStudy (shane_b):

Problems like this?\[(\frac{?}{x+2})(\frac{x+2}{x-3})=x+3\]

OpenStudy (anonymous):

take the unknown as a variable, say suppose y. then simplify yourexpression and get the value of y by solving

OpenStudy (anonymous):

@shane_b yes

OpenStudy (shane_b):

The math behind these is essentially same as any other fractions. In your example, you can immediately cancel out the (x+2)'s on the left side to simplify it.

OpenStudy (anonymous):

in your example, y= (x+3)(x-3)

OpenStudy (anonymous):

so your numerator would be x^2-9

OpenStudy (anonymous):

what does ^ mean?

OpenStudy (shane_b):

exponent: \[x^2\]

OpenStudy (shane_b):

Here's it done in pretty text:\[\frac{y}{x-3}=x+3\]\[y=(x+3)(x-3)\]\[y=x^2-9\]

OpenStudy (shane_b):

Just remember, the normal fraction/equation rules apply...no matter how complicated the problem starts off looking.

OpenStudy (anonymous):

thank you so much

OpenStudy (shane_b):

you're welcome. Good luck :)